Information on OCP's direct materials costs for the current month is as follows: actual quantity of direct materials purchased and used, 30,000 pounds; actual cost of direct materials, $84,000; direct materials efficiency variance, $3,000 U; standard quantity of direct materials allowed for the current month's production, 29,000 pounds. Based on the presented information, OCP's direct materials price variance for the month is closest to

a.

$2,800 F

b.

$2,800 U

c.

$6,000 U

d.

$6,000 F
Answer #1
Answer
The Answer Is D : $ 6,000 F
Material Quantity Variance=Standard Price(Standard Quantity -Actual Quantity)
                    -3,000             =Standard Price(29,000-30000)
                     -3,000           =Standard Price(-1000)
              Standard Price   =$3 per unit
Material Price Variance= Actual quantity required (Standard price - actual price)
                                     =30000(3-(84000/30000))
                                     =30000(3-2.8)
                                      =30000*0.2
                                     =$6,000(Favourable)